# DESI EMISSION LINE GALAXIES: CLUSTERING DEPENDENCE ON STELLAR MASS AND [OII] LUMINOSITY

### v2

@@ 5.1 Application of the Model
Fits to the low, intermediate, and high redshift measurements are shown in Figures 4-6, respectively. Solid curves and shaded regions (top-left, top-right, and bottom-left panels) mark the median and $1\sigma$ bounds of the posterior distribution of predicted 2PCFs. Open circles and uncertainties (bottom-right panels) mark the same but for the number densities.
Our fits are affected by the unexpected spurious signal in the projected 2PCFs, particularly at low redshift. In detail, the model accurately predicts the clustering of high stellar mass samples on all scales (with a few exceptions). However, for most low and intermediate stellar mass samples (especially at lower redshift), the model generally underpredicts the clustering on scales above $r_{p}\sim 10$ $h^{-1}$pc and cannot capture the spurious signal indicated by the measurements. With 93 degrees of freedom, the best-fit $\chi^{2}$ values for the low, intermediate, and high redshift fits are 261.7, 99.9, and 97.4, respectively. The large $\chi^{2}$ of the low-redshift fit suggests that the uncertainties in the measurements are underestimated due to systematic errors not captured in the covariance estimation. Averaged over scales above $10$ $h^{-1}$pc, the fractional uncertainty of the 2PCF measurements at low redshift are 1.3 (1.6) times smaller than those at intermediate (high) redshift. These small uncertainties on large scales, in conjunction with stronger systematic discrepancies, yield a greater penalty in $\chi^{2}$ at low redshift. In contrast, the measurements at intermediate and high redshift are well-fit, indicating the model constraints are appropriately estimated. For the low redshift results, rather than scaling the covariance matrix to produce a lower $\chi^{2}$, we proceed with the caveat that the constraints are not as tight as indicated by the confidence intervals.
@@ 5.2 Constraints on Model Parameters
With constraints placed on the model parameters, we present overall relations to halo mass or stellar mass (according to Section 4) in Figures 7-9 for the low, intermediate, and high redshift samples, respectively. Constraints on all model parameters are given in Appendix C. We first discuss the low-redshift, central galaxy constraints in Figure 7 (thickly outlined panels).
Figure 7: Model constraints from the low-redshift ($1.00.8<z<$ 1.0) samples. From left to right, the model relations are: (top row) occupation fraction of central ELGs, median $\log M_{*}$ of centrals, scatter in $\log M_{*}$ of centrals, (middle row) median $\log L_{\rm[OII]}$ of centrals, scatter in $\log L_{\rm[OII]}$ of centrals, normalization for the conditional stellar mass function of satellites, (bottom row) low-mass slope for the conditional stellar mass function of satellites, median $\log L_{\rm[OII]}$ of satellites, and scatter in $\log L_{\rm[OII]}$ of satellites. Solid lines indicate the median solution. Dark and light shaded regions mark the $1\sigma$ and $2\sigma$ bounds, respectively. Panels for $\mu_{x}$, $\mu_{y}$, and $\mu_{y,\>\rm sat}$ also include gray shaded regions that mark the central 95% range spanned by the sample data. Panels for the constraints on central galaxies have thicker outlines.
Figure 8: Same as Figure 7 but for the $1.21.0<z<$ 1.2 intermediate-redshift samples. For comparison, the median results from the low-redshift samples (Figure 7) are shown as blue dashed lines.
Figure 9: Same as Figure 8 but for the $1.61.2<z<$ 1.6 high-redshift samples.
@@
The median [OII] luminosity of centrals ($\mu_{y}$) increases at low halo mass, flattens within the [OII] luminosity range spanned by our samples (gray band), and decreases at high halo mass. We performed several tests to verify that the model prefers this shape over a flat or monotonically increasing function, which might be expected if [OII] luminosity traces the star formation main sequence. The parabolic behavior is driven by the interplay between $\mu_{y}$ and $A_{c}$. With $\mu_{y}$ steeply decreasing toward low and high halo mass where $A_{c}$ remains non-zero, galaxies in our sample are placed into halos that span a small range of mass. Regardless of how the halo mass selection is performed, the end result remains the same: the model prefers a narrow range of halo masses to produce central ELGs in our samples. The scatter in the $\log L_{\rm[OII]}$ of centrals ($\sigma_{y}$) is $\sim$0.1 dex, consistent with the small measurement uncertainty.
Moving to the satellite constraints (thinly-outlined panels), the amplitude of the satellite distribution ($\phi_{s}$) increases with halo mass but is small, meaning the fraction of galaxies in our samples parameterized as satellites is low. This constraint reflects the relatively flat small-scale clustering in our 2PCF measurements and indicates a weak one-halo term. Due to the low satellite fraction, all other satellite parameters ($\alpha_{s}$, $\mu_{y,\>\rm sat}$, and $\sigma_{y,\>\rm sat}$) are poorly constrained.
@@
At high redshift (Figure 9), $A_{c}$ and $\mu_{y}$ behave differently. Particularly, $A_{c}$ is small and decreases with halo mass, while $\mu_{y}$ remains flat at high halo mass instead of turning over. The change in these two relations highlights their degeneracy as described above. At high redshift, $A_{c}$ is used to exclude high-mass halos, allowing $\mu_{y}$ to remain flat and within the [OII] luminosity range of our samples. At each redshift, either combination of behaviors can be imposed with little penalty in $\chi^{2}$. This degeneracy is further illustrated in Section 5.3, where consequences in the occupation statistics are discussed.
@@ 5.3 Implications on the Occupation Statistics
With constraints placed on the model, Equations 8 and 16 are integrated over each sample’s bounds to compute the mean occupation functions of centrals and satellites, respectively. We show the occupation statistics of each sample as functions of halo mass in Figure 10. Our samples contain a low fraction of satellites, as discussed in Section 5.2 and reinforced by the low satellite occupation functions (dashed lines) in Figure 10.
@@
[…]
At all redshifts, a double-peak behavior that softens with increasing [OII] luminosity is observed, which is a direct consequence of the turnover in the mean [OII] luminosity-halo mass relation ($\mu_{y}$; see Section 5.2). Given the [OII] luminosity range of the lowest [OII] luminosity samples, $\mu_{y}$ crosses the range at two distinct halo mass scales, causing a doubly-peaked occupation function. The non-zero values of $\sigma_{y}$ serve to smooth the distribution. A similar result occurs for the intermediate [OII] luminosity samples, but the two halo mass scales are closer together and the peaks in the occupation functions are less separated. In the case of high [OII] luminosity samples, $\mu_{y}$ approaches but never exceeds the samples’ [OII] luminosity range, creating a singly-peaked occupation function with a maximum at the turnover scale. This behavior is less evident at high redshift because $\mu_{y}$ flattens at higher halo mass while $A_{c}$ suppresses the number of galaxies in high-mass halos.
[…]
Figure 11: Results from an alternative parameterization of the $A_{c}$-$M_{h}$ relation. For illustrative purposes, only constraints for the intermediate redshift samples are shown. Left panels show the occupation fraction of central ELGs (top) and median $\log L_{\rm[OII]}$ of centrals (bottom), formatted as in Figure 7. The right panel shows the mean occupation functions of the intermediate stellar mass samples, formatted as in Figure 10. All other constraints are largely unchanged.
Additional property dependence in the mean occupation functions are evident at high redshift. The mean halo mass increases with [OII] luminosity, indicated by the changing shapes of the occupation functions. There are no trends with stellar mass, which is likely due to the large measurement scatter captured by the model. Theoretically, more precise estimates of stellar mass could reveal trends that are currently unobservable. Inspection of the derived quantities in Section 5.4 further support the presence of galaxy property dependence in the high redshift samples.
@@ 5.4 Implications on the Derived Quantities
Figure 12: Derived quantities of each sample. From top to bottom, the rows are galaxy bias, median halo mass of centrals, satellite fraction, and median halo mass of satellites. Halo masses are in units of $h^{-1}M_{\odot}$. Each panel contains all luminosity and redshift samples at fixed stellar mass. Data points are colored according to the binning in stellar mass and [OII] luminosity in Figure 3. From low to high, redshifts are marked as filled circles, filled stars, and open squares.
@@
Galaxy bias (first row in Figure 12) is tightly constrained and generally increases with stellar mass, especially at high redshift. At lower redshifts, the correlation with stellar mass exists but is weaker. Trends with [OII] luminosity are more redshift-dependent. At low and intermediate redshift, galaxy bias is nearly constant with [OII] luminosity. However, at high redshift, galaxy bias shows an increase with [OII] luminosity, with the highest stellar mass sample having lower significance. To quantify this trend, we perform a linear fit to each series of three [OII] luminosity samples at fixed stellar mass and redshift. All cases at low and intermediate redshift have best-fitting slopes with an absolute value less than 0.04 dex-1. In contrast, the best-fitting slopes at high redshift are much larger: $\left(0.72\pm 0.15\right)$ dex-1, $\left(0.73\pm 0.21\right)$ dex-1, and $\left(0.43\pm 0.28\right)$ dex-1 for the low, intermediate, and high stellar mass samples, respectively. The apparent [OII] luminosity dependence at only high redshift might instead be a consequence of our redshift binning; above $\log\left[L_{\rm{[OII]}}\,/\,\left(\rm{erg\;\,s}^{-1}\right)\right]\sim 42.2$, galaxy bias also increases at lower redshifts, but this trend is most evident at high redshift due to the overall higher [OII] luminosities in that bin. Finally, we find that at fixed stellar mass and [OII] luminosity, galaxy bias increases with redshift.
At all redshifts, the median halo mass of centrals (second row in Figure 12) shows an increase with stellar mass that softens at high [OII] luminosity. Central halo mass also increases with [OII] luminosity, but the prevalence is redshift-dependent. At low and intermediate redshift, the trend flattens at high stellar mass, while the high-redshift case remains strong. At fixed stellar mass and [OII] luminosity, central halo mass decreases with redshift.
@@ 5.5 Comparison With Other Studies
[…]
@@
The central occupation functions produced by our model (Figure 10) also span a small range of halo mass. As discussed in Section 5.3, the double-peak shape observed in several of our occupation functions is insignificant, and the single-peak profile from alternative parameterizations (Figure 11) is similar to that in Prada et al. (2023). The more meaningful quantity to compare is the overall halo mass scale. Our central ELGs consistently occupy higher-mass halos than those in Prada et al. (2023). We find that this discrepancy is largely due to differences in the satellite fraction. Our model classifies nearly all galaxies as centrals; in contrast, the satellite fraction in Prada et al. (2023) is $\sim$13%. With a larger satellite fraction, the central occupation function must shift to lower halo masses to compensate for the higher bias induced by the satellites. Using the halo bias in Tinker et al. (2005), we estimate that with the total galaxy bias fixed, decreasing the satellite fraction in Prada et al. (2023) to 0% will shift the peak of their central occupation function to $\sim$1012.2 $h^{-1}M_{\odot}$, which agrees more with our halo mass scale. Additionally, our sample cuts remove $\sim$10of ELGs with the lowest [OII] luminosities and stellar masses, which might increase the halo mass scale in our model.
@@
Prada et al. (2023) also report, for the entire ELG dataset, galaxy biases that are lower than those of our samples defined by galaxy property (Figure 12). As a test, we perform the clustering measurement on the combination of our nine samples (at fixed redshift) and find a lower bias that is more consistent with Prada et al. (2023). For example, the projected 2PCF amplitude of the combined low-redshift samples at $r_{p}\sim 30$ $h^{-1}$pc is, on average, $\sim$1.9 times lower than those of the individual samples. The excess clustering of our ELG samples on such large scales indicates the likely presence of systematic errors that requires further investigation (e.g., by studying the cross correlation between samples). On smaller scales, the clustering is less affected by such systematic errors, and we can still obtain meaningful constraints on the model.
[…]
@@
Gao et al. (2023) model the central stellar-to-halo mass relation as a double power law and constrain the relation by fitting the stellar mass-dependent clustering of ELGs and luminous red galaxies. The stellar masses used in their work are derived by fitting the DESI broadband photometry with CIGALE (Boquien et al., 2019), which produces stellar masses with an overall offset and smaller dispersion than those used here. In addition, the extent of the offset depends on stellar mass at the high stellar mass end. Despite parameterizing the stellar-to-halo mass relation with a shape similar to Gao et al. (2023), we find that these differences in stellar mass make it difficult to directly compare constraints. Therefore, our stellar-to-halo mass relation should be interpreted specifically for DESI ELGs with stellar masses derived via FastSpecFit.
@@
Many works have explored conformity and its role in the small scale clustering of the DESI ELGs. Yuan et al. (2025) present a physical argument for conformity, in which ELGs have preferentially undergone merger activity, resulting in perturbed morphology and boosted star formation. Such tidal force-driven star formation increases the likelihood that satellites near ELGs are also ELGs, enhancing the clustering on scales of a few hunderd kiloparsecs. Rocher et al. (2023) find that in order to accurately fit the strong small-scale clustering of DESI ELGs below $r_{p}\sim 0.1$ $h^{-1}$pc, conformity is required in the model. Similarly, Gao et al. (2024) report an improved fit to DESI ELG clustering below $r_{p}\sim 0.3$ $h^{-1}$pc when including conformity. Because we divide the data by galaxy property, the clustering on such small scales is not well-measured, limiting our ability to probe conformity. Larger datasets from DESI will enable us to constrain conformity and its possible dependence on galaxy property. Our constraints and interpretation of satellite fractions may also be sensitive to the inclusion of conformity (or lack thereof) as discussed in other studies (Rocher et al., 2023; Ortega-Martinez et al., 2024; Yuan et al., 2025).

### v3

@@ 5.1 Application of the Model
Fits to the low-, intermediate-, and high-redshift measurements are shown in Figures 4-6, respectively. Solid curves and shaded regions (top-left, top-right, and bottom-left panels) mark the median and $1\sigma$ bounds of the posterior distribution of predicted 2PCFs. Open circles and uncertainties (bottom-right panels) mark the same but for the number densities.
Our goodness of fits are affected by the unexpected spurious signal in the projected 2PCFs, particularly at low redshift. In detail, the model accurately predicts the clustering of high-stellar-mass samples on all scales (with a few exceptions). However, for most low- and intermediate-stellar-mass samples (especially at lower redshift), the model generally underpredicts the clustering on scales above $r_{p}\sim 10$ $h^{-1}$Mpc and cannot capture the spurious signal indicated by the measurements. With 93 degrees of freedom, the best-fit $\chi^{2}$ values for the low-, intermediate-, and high-redshift fits are 261.7, 99.9, and 97.4, respectively. The large $\chi^{2}$ of the low-redshift fit suggests that the uncertainties in the measurements are underestimated due to systematic errors not captured in the covariance estimation. Averaged over scales above 1010 $h^{-1}$Mpc, the fractional uncertainty of the 2PCF measurements at low redshift are 1.3 (1.6) times smaller than those at intermediate (high) redshift. These small uncertainties on large scales, in conjunction with stronger systematic discrepancies, yield a greater penalty in $\chi^{2}$ at low redshift. In contrast, the measurements at intermediate and high redshift are well fit, indicating the model constraints are appropriately estimated.
[…]
@@ 5.2 Constraints on Model Parameters
With constraints placed on the model parameters, we present overall relations to halo mass or stellar mass (according to Section 4) in Figures 7-9 for the low-, intermediate-, and high-redshift samples, respectively. Constraints on all model parameters are given in Appendix C. We first discuss the low-redshift central galaxy constraints in Figure 7 (thickly outlined panels).
Figure 7: Model constraints from the low-redshift ($0.8<z<1.0$) samples. From left to right, the model relations are (top row) the occupation fraction of central ELGs, the median $\log M_{*}$ of centrals, the scatter in $\log M_{*}$ of centrals, (middle row) the median $\log L_{\rm[OII]}$ of centrals, the scatter in $\log L_{\rm[OII]}$ of centrals, the normalization for the conditional stellar mass function of satellites, (bottom row) the low-mass slope for the conditional stellar mass function of satellites, the median $\log L_{\rm[OII]}$ of satellites, and the scatter in $\log L_{\rm[OII]}$ of satellites. Solid lines indicate the median solution. Dark and light shaded regions mark the $1\sigma$ and $2\sigma$ bounds, respectively. Panels for $\mu_{x}$, $\mu_{y}$, and $\mu_{y,\>\rm sat}$ also include gray shaded regions that mark the central 95% range spanned by the sample data. Panels for the constraints on central galaxies have thicker outlines.
Figure 8: Same as Figure 7 but for the $1.0<z<1.2$ intermediate-redshift samples. For comparison, the median results from the low-redshift samples (Figure 7) are shown as blue dashed lines.
Figure 9: Same as Figure 8 but for the $1.2<z<1.6$ high-redshift samples.
@@
The median [OII] luminosity of centrals ($\mu_{y}$) increases at low halo mass, flattens within the [OII] luminosity range spanned by our samples (gray band), and decreases at high halo mass. We performed several tests to verify that the model prefers this shape over a flat or monotonically increasing function, which might be expected if [OII] luminosity traces the star formation main sequence. The parabolic behavior is driven by the interplay between $\mu_{y}$ and $A_{c}$. With $\mu_{y}$ steeply decreasing toward low and high halo mass where $A_{c}$ remains nonzero, galaxies in our sample are placed into halos that span a small range of mass. Regardless of how the halo mass selection is performed, the end result remains the same: The model prefers a narrow range of halo masses to produce central ELGs in our samples. The scatter in the $\log L_{\rm[OII]}$ of centrals ($\sigma_{y}$) is $\sim$0.1 dex, consistent with the small measurement uncertainty.
Moving to the satellite constraints (thinly outlined panels), the amplitude of the satellite distribution ($\phi_{s}$) increases with halo mass but is small, meaning the fraction of galaxies in our samples parameterized as satellites is low. This constraint reflects the relatively flat small-scale clustering in our 2PCF measurements and indicates a weak one-halo term. Due to the low satellite fraction, all other satellite parameters ($\alpha_{s}$, $\mu_{y,\>\rm sat}$, and $\sigma_{y,\>\rm sat}$) are poorly constrained.
@@
At high redshift (Figure 9), $A_{c}$ and $\mu_{y}$ behave differently. In particular, $A_{c}$ is small and decreases with halo mass, while $\mu_{y}$ remains flat at high halo mass instead of turning over. The change in these two relations highlights their degeneracy as described above. At high redshift, $A_{c}$ is used to exclude high-mass halos, allowing $\mu_{y}$ to remain flat and within the [OII] luminosity range of our samples. At each redshift, either combination of behaviors can be imposed with little penalty in $\chi^{2}$. This degeneracy is further illustrated in Section 5.3, where consequences in the occupation statistics are discussed.
@@ 5.3 Implications on the Occupation Statistics
With constraints placed on the model, Equations (8) and (16) are integrated over each sample’s bounds to compute the mean occupation functions of centrals and satellites, respectively. We show the occupation statistics of each sample as functions of halo mass in Figure 10. Our samples contain a low fraction of satellites, as discussed in Section 5.2 and reinforced by the low satellite occupation functions (dashed lines) in Figure 10.
@@
All central occupation functions (solid lines) at low and intermediate redshift span a narrow range of halo masses, $\sim$$10^{12}$ $h^{-1}M_{\odot}$ to $10^{12.7}$ $h^{-1}M_{\odot}$. At high redshift, the central occupation functions have tails extending to higher halo mass but preserve the same median value by also adding low-mass halos.
At all redshifts, a double-peak behavior that softens with increasing [OII] luminosity is observed, which is a direct consequence of the turnover in the mean [OII] luminosity-halo mass relation ($\mu_{y}$; see Section 5.2). Given the [OII] luminosity range of the lowest [OII] luminosity samples, $\mu_{y}$ crosses the range at two distinct halo mass scales, causing a doubly peaked occupation function. The nonzero values of $\sigma_{y}$ serve to smooth the distribution. A similar result occurs for the intermediate [OII] luminosity samples, but the two halo mass scales are closer together and the peaks in the occupation functions are less separated. In the case of high [OII] luminosity samples, $\mu_{y}$ approaches but never exceeds the samples’ [OII] luminosity range, creating a singly peaked occupation function with a maximum at the turnover scale. This behavior is less evident at high redshift because $\mu_{y}$ flattens at higher halo mass while $A_{c}$ suppresses the number of galaxies in high-mass halos.
As described earlier, the model uses the freedom in $\mu_{y}$ to apply a halo mass selection that cannot be captured by $A_{c}$ due to its monotonicity. To push all selection effects into $A_{c}$, alternative parameterizations with varying flexibility were tested. One example is shown in Figure 11, where $A_{c}$ varies with halo mass similar to $\mu_{y}$, providing a means to asymmetrically decrease at both low and high halo mass. The model has minimal preference toward the original parameterization in Figure 8 ($\chi^{2}\sim 99.9$ with 93 degrees of freedom) over that in Figure 11 ($\chi^{2}\sim 101.0$ with 92 degrees of freedom), indicating that the double-peak feature in the occupation function is insignificant. In all cases, the data push the model toward a scenario in which DESI preferentially observes ELGs within a narrow range of halo masses centered on $\sim$$10^{12.4}$ $h^{-1}M_{\odot}$. This preference may be a consequence of the grzgrz selection used to define DESI ELG targets (Raichoor et al., 2023). By imposing a faint cut on the $g$-band fiber magnitude of $g_{\rm fib}<24.1$, faint ELGs in low-mass halos lack enough UV luminosity to be targeted by DESI. Similarly, by requiring blue colors in grg-r and rzr-z, quiescent galaxies that dominate high-mass halos do not enter the selection.
Figure 11: Results from an alternative parameterization of the $A_{c}$-$M_{h}$ relation. For illustrative purposes, only constraints for the intermediate-redshift samples are shown. Left panels show the occupation fraction of central ELGs (top) and median $\log L_{\rm[OII]}$ of centrals (bottom), formatted as in Figure 7. The right panel shows the mean occupation functions of the intermediate-stellar-mass samples, formatted as in Figure 10. All other constraints are largely unchanged.
Additional property dependence in the mean occupation functions is evident at high redshift. The mean halo mass increases with [OII] luminosity, indicated by the changing shapes of the occupation functions. There are no trends with stellar mass, which is likely due to the large measurement scatter captured by the model. Theoretically, more precise estimates of stellar mass could reveal trends that are currently unobservable. Inspection of the derived quantities in Section 5.4 further supports the presence of galaxy property dependence in the high-redshift samples.
@@ 5.4 Implications on the Derived Quantities
Figure 12: Derived quantities of each sample. From top to bottom, the rows are the galaxy bias, the median halo mass of centrals, the satellite fraction, and the median halo mass of satellites. Halo masses are in units of $h^{-1}M_{\odot}$. Each panel contains all luminosity and redshift samples at fixed stellar mass. Data points are colored according to the binning in stellar mass and [OII] luminosity in Figure 3. From low to high, redshifts are marked as filled circles, filled stars, and open squares.
@@
Galaxy bias (first row in Figure 12) is tightly constrained and generally increases with stellar mass, especially at high redshift. At lower redshifts, the correlation with stellar mass exists but is weaker. Trends with [OII] luminosity are more redshift dependent. At low and intermediate redshift, galaxy bias is nearly constant with [OII] luminosity. However, at high redshift, galaxy bias shows an increase with [OII] luminosity, with the highest-stellar-mass sample having lower significance. To quantify this trend, we perform a linear fit to the galaxy bias as a function of (logarithmic) [OII] luminosity at fixed stellar mass and redshift (i.e., three data points are used in each linear fit). All cases at low and intermediate redshift have best-fitting slopes with an absolute value less than 0.04 dex-1. In contrast, the best-fitting slopes at high redshift are much larger: $\left(0.72\pm 0.15\right)$ dex-1, $\left(0.73\pm 0.21\right)$ dex-1, and $\left(0.43\pm 0.28\right)$ dex-1 for the low-, intermediate-, and high-stellar-mass samples, respectively. The apparent [OII] luminosity dependence at only high redshift might instead be a consequence of our redshift binning; above $\log\left[L_{\rm{[OII]}}\,/\,\left(\rm{erg\;\,s}^{-1}\right)\right]\sim 42.2$, galaxy bias also increases at lower redshifts, but this trend is most evident at high redshift due to the overall higher [OII] luminosities in that bin. Finally, we find that at fixed stellar mass and [OII] luminosity, galaxy bias increases with redshift.
At all redshifts, the median halo mass of centrals (second row in Figure 12) shows an increase with stellar mass that softens at high [OII] luminosity. Central halo mass also increases with [OII] luminosity, but the prevalence is redshift dependent. At low and intermediate redshift, the trend flattens at high stellar mass, while the high-redshift case remains strong. At fixed stellar mass and [OII] luminosity, central halo mass decreases with redshift.
@@ 5.5 Comparison With Other Studies
Prada et al. (2025) use a modified SHAM approach to place DESI ELGs in halos and produce mock lightcones. Their mean halo occupation function for centrals resembles a Gaussian that peaks at $\sim$1011.7 $h^{-1}M_{\odot}$, and that for the satellites increases like a power law, approaching unity at $\sim$$10^{14.6}$ $h^{-1}M_{\odot}$. Both shapes are consistent with those predicted by semi-analytical models in Gonzalez-Perez et al. (2020).
@@
The central occupation functions produced by our model (Figure 10) also span a small range of halo mass. As discussed in Section 5.3, the double-peak shape observed in several of our occupation functions is insignificant, and the single-peak profile from alternative parameterizations (Figure 11) is similar to that in Prada et al. (2025). The more meaningful quantity to compare is the overall halo mass scale. Our central ELGs consistently occupy higher-mass halos than those in Prada et al. (2025). We find that this discrepancy is largely due to differences in the satellite fraction. Our model classifies nearly all galaxies as centrals; in contrast, the satellite fraction in Prada et al. (2025) is $\sim$13%. With a larger satellite fraction, the central occupation function must shift to lower halo masses to compensate for the higher bias induced by the satellites. Using the halo bias in Tinker et al. (2005), we estimate that with the total galaxy bias fixed, decreasing the satellite fraction in Prada et al. (2025) to 0% will shift the peak of their central occupation function to $\sim$1012.2 $h^{-1}M_{\odot}$, which agrees more with our halo mass scale. Additionally, our sample cuts remove $\sim$10% of ELGs with the lowest [OII] luminosities and stellar masses, which might increase the halo mass scale in our model.
@@
Prada et al. (2025) also report, for the entire ELG data set, galaxy biases that are lower than those of our samples defined by galaxy property (Figure 12). As a test, we performed the clustering measurement on the combination of our nine samples (at fixed redshift) and found a lower bias that is more consistent with Prada et al. (2025). For example, the projected 2PCF amplitude of the combined low-redshift samples at $r_{p}\sim 30$ $h^{-1}$Mpc is, on average, $\sim$1.9 times lower than those of the individual samples. The excess clustering of our ELG samples on such large scales indicates the likely presence of systematic errors that requires further investigation (e.g., by studying the cross correlation between samples). On smaller scales, the clustering is less affected by such systematic errors, and we can still obtain meaningful constraints on the model.
Regarding satellite fractions, this study is not unique in finding low values for ELGs. Yu et al. (2024) apply a SHAM method to ELGs in the DESI One-Percent Survey and report satellite fractions below $6\%$ and consistent with zero at the $2.5\sigma$ level. Beyond DESI, HOD modeling of galaxies with the highest [OII] luminosity in the VIMOS Public Extragalactic Redshift Survey (Scodeggio et al., 2018) has revealed satellite fractions of $\sim$$5\%\pm 2\%$ (Gao et al., 2022). While our result is in agreement with a satellite fraction as low as zero, many samples have loose constraints that are also consistent with up to $\sim$$10-20\%$ as found in other studies (e.g., Lin et al. 2023; Prada et al. 2025). For a proper comparison with any of these studies, the modeling details and sample definitions must be closely examined.
@@
Gao et al. (2023) model the central SHMR as a double power law and constrain the relation by fitting the stellar-mass-dependent clustering of ELGs and luminous red galaxies. The stellar masses used in their work are derived by fitting the DESI broadband photometry with CIGALE (Boquien et al., 2019), which produces stellar masses with an overall offset and smaller dispersion than those used here. In addition, the extent of the offset depends on stellar mass at the high-stellar-mass end. Despite parameterizing the SHMR with a shape similar to Gao et al. (2023), we find that these differences in stellar mass make it difficult to directly compare constraints. Therefore, our SHMR should be interpreted specifically for DESI ELGs with stellar masses derived via FastSpecFit.
@@
Galaxy conformity, defined as the correlation of properties between neighboring galaxies such as centrals and satellites (Weinmann et al., 2006), and its role in the small-scale clustering of the DESI ELGs have been explored in several studies. Yuan et al. (2025) present a physical argument for conformity, in which ELGs have preferentially undergone merger activity, resulting in perturbed morphology and boosted star formation. Such tidal-force-driven star formation increases the likelihood that satellites near ELGs are also ELGs, enhancing the clustering on scales of a few hundred kiloparsecs. Rocher et al. (2023) find that in order to accurately fit the strong small-scale clustering of DESI ELGs below $r_{p}\sim 0.1$ $h^{-1}$Mpc, conformity is required in the model. Similarly, Gao et al. (2024) report an improved fit to DESI ELG clustering below $r_{p}\sim 0.3$ $h^{-1}$Mpc when including conformity. Because we divide the data by galaxy property, the clustering on such small scales is not well measured, limiting our ability to probe conformity. Larger data sets from DESI will enable us to constrain conformity and its possible dependence on galaxy property. Our constraints and interpretation of satellite fractions may also be sensitive to the inclusion of conformity (or lack thereof) as discussed in other studies (Rocher et al., 2023; Ortega-Martinez et al., 2024; Yuan et al., 2025).